Expand and simplify: (8 + 2)2 O 64 +22 O 2 + 162 +16 O 2 +167 + 64 Oz+16
Answer #1

(8+Z)~ = 189+2.8.2+22 = (64+16 2 + zy 3rd aftion is connect.